# Predicting Bad Goods Risk Scores with ARIMA Time Series: A Novel Risk Assessment Approach

### v2

@@ Abstract
Predictive Analytics, Bad Goods, Time Series ARIMA, Risk Assessment, Quality Control, Supply Chain Management, Defect Prediction
@@ I INTRODUCTION
In modern industrial systems, detecting and preventing defective or substandard products—termed “bad goods”—such as manufacturing flaws or spoiled items like Organic Beer-G 1 Liter, remains a critical challenge. These defects result in financial losses, reputational harm, and supply chain inefficiencies. Traditional approaches like statistical process control and manual inspections struggle to address the complexity of large-scale operations [1]. The advent of big data and advanced analytics has elevated predictive methods as a key strategy for preempting such risks [2].
This study introduces a predictive analytics framework that integrates Time Series ARIMA models with risk assessment to forecast bad goods occurrences and guide supply chain interventions. Leveraging time series data—encompassing sales, returns, and capacity—ARIMA effectively models temporal patterns, outperforming simpler methods like Exponential Smoothing in capturing linear trends [2]. A risk assessment layer, inspired by Failure Mode and Effects Analysis (FMEA), quantifies failure likelihood and impact, enhancing decision-making [3].
Motivated by the limitations of reactive quality management, this research aims to deliver a proactive, data-driven solution for industries such as beverage production and retail. The objectives are to: (1) build a precise ARIMA-based predictive model, (2) incorporate risk assessment for practical insights, and (3) benchmark its efficacy against existing standards. By reducing bad goods, this approach promises cost savings and improved customer satisfaction across manufacturing, retail, and logistics sectors.
@@ II BASIC CONCEPTS
Time Series Analysis (ARIMA): Time series analysis involves examining data points collected sequentially over time to identify trends, seasonality, and anomalies. In the context of bad goods, time series data may include production metrics, sales, returns, and capacity, exhibiting temporal patterns indicative of quality issues. The Autoregressive Integrated Moving Average (ARIMA) model, a classical time series forecasting technique, is widely applied to predict such data by modeling linear relationships and stationarity [1]. ARIMA’s strengths lie in its interpretability and effectiveness for stationary or transformed non-stationary data, although it may struggle with highly non-linear or complex patterns compared to advanced machine learning methods [2].
Risk Assessment: Risk assessment quantifies the likelihood and impact of adverse events, enabling prioritized decision-making in supply chains. In quality management, techniques such as Failure Mode and Effects Analysis (FMEA) evaluate the severity, occurrence, and detectability of defects, providing a structured approach to risk scoring [3]. By integrating ARIMA-based predictions with risk scoring, this research extends beyond forecasting to offer actionable insights, addressing the proactive needs of industries facing bad goods risks. While traditional risk models often operate independently of predictive analytics, their synergy in this framework bridges a critical gap.
Quality Control: Quality control in supply chains ensures product integrity, minimizing defects and returns. For perishable goods like beer, factors such as freshness, shelf life, and capacity play a pivotal role in determining bad goods risks. This study leverages quality control principles to guide ARIMA predictions and risk assessments, ensuring alignment with operational realities and industry standards.
Together, these concepts form the backbone of a predictive analytics system tailored to bad goods for Organic Beer-G 1 Liter. Time Series ARIMA provides the forecasting foundation, risk assessment ensures practical applicability, and quality control principles enhance relevance. This study builds on these principles to develop a unified model that advances quality assurance in supply chain settings.
@@ III RELATED WORKS
The application of predictive analytics to quality control and defect detection in supply chains has garnered increasing attention in recent years. Traditional statistical methods, such as Autoregressive Integrated Moving Average (ARIMA) models, have been extensively utilized for time series forecasting in industrial contexts [1]. While effective for linear and stationary data, ARIMA models face challenges with non-linear, high-dimensional, or noisy datasets common in modern supply chains, prompting exploration of alternative approaches [2]. For instance, machine learning techniques like Support Vector Machines (SVM) and Random Forests have been applied for anomaly detection and defect prediction in manufacturing processes, but these methods often struggle to model long-term temporal dependencies critical for bad goods prediction.
Advanced time series methods, particularly deep learning approaches, have revolutionized forecasting in supply chains. Recurrent Neural Networks (RNNs), especially Long Short-Term Memory (LSTM) networks, have proven effective for modeling sequential data and predicting equipment failures or product defects over extended periods [4]. Convolutional Neural Networks (CNNs) have also been adapted for time series tasks, extracting local features from multivariate inputs to improve prediction accuracy [5]. Hybrid models combining CNNs and LSTMs have shown promise in supply chain demand forecasting, as demonstrated by [2], yet their computational complexity and data requirements can limit scalability for real-time applications. Despite these advancements, few studies have specifically targeted bad goods prediction, leaving a gap in quality-specific predictive frameworks for supply chains.
@@ IV-A Data Preprocessing Phase
The first phase, depicted in the top section of Figure 1, focuses on collecting and preprocessing historical data from the supply chain. Data originates from the manufacturer, which distributes products through multiple distributors to various retailers. The data flow involves:
@@
- Data Collection: The manufacturer sends data to distributors, who then forward it to retailers. This hierarchical flow ensures comprehensive coverage of the supply chain.
@@ IV-C Risk Score Calculation Phase (Bad Goods)
  • Return Rate (%): Forecasted monthly/weekly for the next calendar year, based on historical patterns.
@@
  • Expected Return Quantity:
@@
  • Retailer Inventory Capacity (Monthly): Forecasted and derived based on historical patterns.
  • Freshness Ratio:
@@
  • Bad Good Risk Score:
@@ VI-H Interpretation of Bad Goods Risk Scores
The ‘BG Risk Score‘ in Table I categorizes the risk of Organic Beer-G 1 Liter becoming a bad good (e.g., defective, expired, or returned) into three levels: Low Risk (score $<$ 0.4), Medium Risk (0.4 $\leq$ score $<$0.8), and High Risk (score $\geq$ 0.8). These levels indicate the likelihood of bad goods, with High Risk requiring immediate action, Medium Risk suggesting preventive measures, and Low Risk recommending monitoring. Based on this classification, we analyze the risk scores for each month and propose actionable recommendations to mitigate risks by adjusting ‘Demand Plan (Sales Qty)‘, ‘Retailer Inventory Capacity (Monthly)‘, or ‘Freshness Left (Months)‘:

### v3

@@ Abstract
Predictive Analytics, Bad Goods, Time Series ARIMA, Risk Assessment, Quality Control, Supply Chain Management.
@@ I INTRODUCTION
In modern industrial systems, identifying and mitigating defective, expired or substandard products, called “bad goods”—such as manufacturing defects or unsold perishable items like Organic Beer-G 1 Liter poses a significant challenge. These issues lead to financial losses, reputational damage, and supply chain disruptions. Conventional methods, such as statistical process control and manual planning, are often inadequate to manage the complexity of large-scale operations [1]. The rise of big data and advanced analytics has placed predictive approaches as a vital tool for proactively addressing these risks [2].
This research presents a predictive analytics framework that combines Time Series ARIMA models with risk assessment to anticipate occurrence of bad goods and inform supply chain strategies. Using time-series data such as sales, returns, and production capacity, ARIMA models effectively capture temporal patterns, surpassing simpler techniques such as Exponential Smoothing in detecting linear trends [2].A risk assessment layer, inspired by Failure Mode and Effects Analysis (FMEA), quantifies failure likelihood and impact, enhancing decision-making [3].
Motivated by the limitations of reactive quality management, this research aims to deliver a data-driven proactive solution for industries such as beverage production and retail. The objectives are to: (1) build a precise ARIMA-based predictive model, (2) incorporate risk assessment for practical insights, and (3) benchmark its efficacy against existing standards. By reducing bad goods, this approach promises cost savings and improved customer satisfaction across manufacturing, retail, and logistics sectors.
@@ II BASIC CONCEPTS
Time-Series Analysis (ARIMA): Time-series analysis involves examining data points collected sequentially over time to identify trends, seasonality, and anomalies. In the context of bad goods, time series data may include production metrics, sales, returns, and capacity, exhibiting temporal patterns indicative of quality issues. The autoregressive integrated moving average model (ARIMA), a well-established time series forecasting method, is extensively used to predict data by capturing linear relationships and ensuring the direct separation of stationarity [1]. ARIMA excels in its interpretability and suitability for stationary or differenced non-stationary data, though it may falter with highly non-linear or intricate patterns compared to advanced machine learning approaches [2].
Risk Assessment: Risk assessment assesses the risk likelihood and impacts in order to support prioritization of supply chain decision. In quality management, methodologies like the Failure Mode and Effects Analysis (FMEA) score defect severity, occurrence, and detectability, providing an organized approach to risk ranking [3]. The integration of risk scoring with ARIMA-based forecasts in this study provides a forward leap from conventional forecasting, offering proactive industry solutions to bad goods risk management. Unlike stand-alone risk models, the integrated model fills an important void by utilizing predictive analytics.
Quality Control: The quality control in the supply chains is the protection of the product integrity and therefore minimization of defects and returns. For perishable products such as beer, the freshness, shelf-life, production capabilities are important in analyzing the risk of bad goods. This study embeds quality control concepts to direct ARIMA forecasting and risk analysis to make certain that they fit to operational realities and industry norms.
Collectively, these elements underpin a predictive analytics system designed for bad goods management in Organic Beer-G 1 Liter. Time-Series ARIMA provides robust forecasting, risk assessment ensures practical relevance, and quality control enhances operational alignment. This study integrates these components to develop a cohesive model that advances quality assurance in supply chain environments.
@@ III RELATED WORKS
Using predictive analytics in quality control and defect detection for supply chain operations is a growing research area. Traditional time series forecasting methods, such as the Autoregressive Integrated Moving Average (ARIMA) model, have been widely used for time series forecasting in industrial applications [1]. However, ARIMA is designed for linear and stationary data, and it is not suitable for non-linear, high-dimensional, or noisy data which is characteristic of modern supply chains, leading to alternate solution methodologies [2]. Machine learning algorithms, such as Support Vector Machines (SVM) and Random Forests, have been applied to anomaly detection and defect prediction in manufacturing, but they often cannot model long-term temporal dependencies necessary for bad goods prediction [9][11].
Advanced time series methods, particularly deep learning approaches, have revolutionized forecasting in supply chains. Recurrent Neural Networks (RNNs), especially Long Short-Term Memory (LSTM) networks, have proven effective for modeling sequential data and predicting equipment failures or product defects over extended periods [14]. Convolutional Neural Networks (CNNs) have also been adapted for time series tasks, extracting local features from multivariate inputs to improve prediction accuracy [15]. Hybrid models combining CNNs, LSTMs and BiLSTMs have shown promise in supply chain demand forecasting, as demonstrated by [2][8], yet their computational complexity and data requirements can limit scalability for real-time applications. Despite these advancements, few studies have specifically targeted bad goods prediction, leaving a gap in quality-specific predictive frameworks for supply chains.
@@ IV-A Data Preprocessing Phase
The first phase, shown in the top section of Figure 1, focuses on collecting and preprocessing historical data from the supply chain. Data originates from the manufacturer, which distributes products through multiple distributors to various retailers. The data flow involves:
@@
- Data Collection: The data is collected from manufacturer who store data at each level like distributors, who then it to retailers. This hierarchical flow ensures comprehensive coverage of the supply chain.
@@ IV-C Risk Score Calculation Phase (Bad Goods)
  • Return Rate (%): Forecasted monthly for the next calendar year, based on historical patterns. This metric indicates the percentage of sold items that are returned.
@@
  • Expected Return Quantity: Represents the anticipated number of returns based on sales volume and return rate. Useful for inventory planning.
@@
  • Retailer Inventory Capacity (Monthly): Forecasted and derived based on historical patterns. It measures the available inventory space after accounting for reruns.
@@
  • Freshness Ratio: Quantifies the freshness of inventory relative to its shelf life. A higher ratio indicates fresher stock.
[…]
  • Bad Good Risk Score: Assesses the risk of inventory spoilage or obsolescence by comparing expected returns to inventory capacity, adjusted by freshness.
@@ VI-H Interpretation of Bad Goods Risk Scores
The ‘BG Risk Score‘ in Table I categorizes the risk of Organic Beer-G 1 Liter becoming a bad good (e.g., defective, expired, or returned) into three levels: Low Risk (score $<$ 0.4), Medium Risk (0.4 $\leq$ score $<$0.8), and High Risk (score $\geq$ 0.8) calculated using equation 1-5. These levels indicate the likelihood of bad goods, with High Risk requiring immediate action, Medium Risk suggesting preventive measures, and Low Risk recommending monitoring. Based on this classification, we analyze the risk scores for each month and propose actionable recommendations to mitigate risks by adjusting ‘Demand Plan (Sales Qty)‘, ‘Retailer Inventory Capacity (Monthly)‘, or ‘Freshness Left (Months)‘: